Alocal group claims that the police issue at least 60parking tickets a day in their area. Toprove their point, they randomlyselect two weeks. Their research yields the number of tickets issued foreachday. Thedataarelistedbelow.Ata 0.01,testthegroupsclaim. Use StatCrunchto 5. calculate the relevant statistics. 20, 48, 41, 68, 69, 70, 57, 60, 83, 32, 60, 72, 58, 63
Answer #1

Data:    

n = 14   

μ = 60   

s = 16.917   

x-bar = 57.214   

Hypotheses:    

Ho: μ ≥ 60   

Ha: μ < 60   

Decision Rule:    

α = 0.05   

Degrees of freedom = 14 - 1 = 13

Critical t- score = -1.770933383   

Reject Ho if t < -1.770933383   

Test Statistic:    

SE = s/√n = 16.917/√14 = 4.521258429

t = (x-bar - μ)/SE = (57.214 - 60)/4.52125842943248 = -0.616200123

p- value = 0.274200458   

Decision (in terms of the hypotheses):    

Since -0.616200123 > -1.770933383 we fail to reject Ho

Conclusion (in terms of the problem):    

There is no sufficient evidence that μ < 60. The claim made by the group is valid.